OpenStudy (anonymous):

x^2+x+7>19 (Solve for x)

OpenStudy (anonymous):

factor \[x^2+x-12>0\] \[(x+4)(x-3)>0\]

OpenStudy (anonymous):

now this animal is a parabola that faces up . it is zero at -4 and 3 so you can tell from the picture (in your head) that it is positive from \[(-\infty,-4)\] and on \[(3,\infty)\]

OpenStudy (anonymous):

with the understanding that "positive" is a synonym for "above the x axis"

OpenStudy (amistre64):

we can start it by subtracting 19 from both sides

OpenStudy (amistre64):

x^2+x+7-19>19-19 x^2 +x -12 > 0 ; now its just a quadratic to factor

OpenStudy (amistre64):

what 2 numbers multiply to get 12?

OpenStudy (amistre64):

1*12 2*6 3*4 so we got to see which set gets us a value for the middle# of +1 12 __ 1 = 1 will a + or - make this happen? no 6 ___ 2 = 1 can a + or - make this happen? no 4 __ 3 = 1 ; this is a good bet; use 4-3 (x+4)(x-3)

OpenStudy (amistre64):

now we determine what = 0 x+4 = 0 when x=? x-3 = 0 when x=? and use those value on our number line

OpenStudy (anonymous):

i hate this method!

OpenStudy (amistre64):

<.....-4.......3......> _ + + ; for (x+4) _ _ + ; for (x-3) ------------------ + - + ; multiplied together we get the results

OpenStudy (amistre64):

so our answers is as sat gave: (-inf,-4) U (3,inf)

OpenStudy (anonymous):

although i do like our explanation

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

"your" explanation i meant

OpenStudy (anonymous):

graph please? to make sure i did it right :)

OpenStudy (anonymous):

anybody?